OpenStudy (anonymous):

Calculate the equation of the line between the points (-6,-1) and (2,-1)

OpenStudy (anonymous):

ideas? what is the slope?

OpenStudy (anonymous):

I got 0/8 for the slope. i did y2-y1 and x2-x1

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

slope is 0. But the other question is, where does it hit the y-axis?

OpenStudy (anonymous):

what do i do after that??

OpenStudy (anonymous):

Let's do this, one equation for a line is y = mx + b Do you know that one?

OpenStudy (anonymous):

I thought I use y-y=m(x-x)

OpenStudy (anonymous):

OK. We just found m. Remember what it is?

OpenStudy (anonymous):

ya But the slope 0/8 confuses me

OpenStudy (anonymous):

0/8 is just 0 forget about the 8

OpenStudy (anonymous):

ok so y= mx+b is the equation?

OpenStudy (anonymous):

We can stick with yours. Since m = 0,the right hand side is 0

OpenStudy (anonymous):

Now we have that y - y0 = 0 And we know one y that is on the line, that is, one y0...

OpenStudy (anonymous):

would the x= -1 for the answer?

OpenStudy (anonymous):

Almost, we had y - y0 = 0 Now we have y0 = -1 y = ?

OpenStudy (anonymous):

y=-1

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

And it makes sense, sketch the points and you'll see.

OpenStudy (anonymous):

hahah thanks alot!! :)

OpenStudy (anonymous):

yw